# NOBLE – Neural Operator with Biologically-informed Latent Embeddings to Capture Experimental Variability in Biological Neuron Models

### v3

@@ 1 Introduction
The challenges of scalability and the computational cost associated with traditional numerical modeling approaches, such as numerical integrators and evolutionary optimization algorithms, have led the scientific community to explore the use of machine learning to accelerate simulations by learning underlying relationships between variables directly from experimental and synthetic data. While neural networks have been used successfully for many applications, they learn mappings between finite-dimensional vectors, which can limit their ability to model physical phenomena that are better described using functions in infinite-dimensional spaces and functional relationships between them [25]. As a result, neural networks can overfit to the training discretization and suffer from limited out-of-distribution capabilities. Neural operators [25, 26, 27] are a principled way to generalize neural networks to learn operators mapping functions to functions, with a universal operator approximation property [28]. A variety of neural operators have been proposed, such as the Fourier Neural Operator (FNO) [29].
@@ 3.3 Neural Operators for Neuron Dynamics Simulation
We choose to use neural operators as they offer clear advantages for modeling complex dynamics (Section˜C.3). Among neural operators, the Fourier Neural Operator (FNO) is very efficient as it leverages fast Fourier transforms on equidistant grids, which aligns naturally with our setting where both experimental recordings and PDE simulations are sampled at constant timesteps. FNOs provide a principled and efficient framework for modeling neuronal dynamics by learning mappings from input currents to voltage responses across a broad family of neuron models and current injections. Unlike conventional neural networks that operate on vector inputs and outputs of fixed sizes, the FNO learns operators, that is, mappings between functions. By operating in the frequency domain, the FNO efficiently captures global, nonlinear, and high-frequency components of voltage responses. These properties allow the model to generalize across different temporal resolutions, input currents, and neuron types, enabling the accurate simulation of unseen configurations without retraining.
@@ 4.1 Experimental Setup
For evaluating NOBLE, we focus on the PVALB neuron example introduced in Section˜2, and additionally validate generality of the framework on a VIP neuron. Each neuron model $\text{HoF}_{\ell}$ receives as input the stacked embeddings of the normalized model features $(I_{\textit{thr}},s_{\textit{thr}})$ and the applied current injection $I$, and outputs the somatic voltage response. The current injections are square DC steps with an activation duration of $400\mathrm{ms}$, consistent across all stimuli used for training and testing. Given a PVALB model $\text{HoF}_{\ell}$ and a current injection $I$, the neural operator in NOBLE takes the stacked embeddings of the (normalized) model features $I_{\textit{thr}}$ and $s_{\textit{thr}}$ associated to $\text{HoF}_{\ell}$ (with $K=1$ frequency) and $I$ (with $K=9$ different frequencies), and outputs the somatic voltage response. We have access to 60 HoF models, where 50 are used during training, $\{\text{HoF}^{\textit{train}}\}$, and the remaining 10 HoF models $\{\text{HoF}^{\textit{test}}\}$ are used for testing. Figure˜10 displays the latent representations in normalized $(I_{\textit{thr}},s_{\textit{thr}})$-space of these HoF models. For more details on the data generation, see Section˜D.1.
@@ 4.2 Testing on HoF Models Included in the Training Set
Figure 3: A) F–I curves from experimental recordings, PDE simulations, and NOBLE predictions on $\{\text{HoF}^{\textit{train}}\}$. For one HoF model, B) compares experimental voltage responses with PDE simulations at a current injection of $0.1\mathrm{nA}$ (top) and $0.11\mathrm{nA}$ (bottom), and C) compares the corresponding PDE simulation with the NOBLE prediction for the same HoF model.
@@ 4.3 Interpolating Between Models
Figure 4: A) F–I curves from experimental recordings, PDE simulations, and NOBLE predictions on 50 interpolated HoF models. For one interpolated model, B) compares experimental voltage responses with PDE simulations at a current injection of $0.1\mathrm{nA}$, and C) compares the corresponding PDE simulation with the NOBLE prediction for the same interpolated HoF model.
We now test NOBLE’s ability to interpolate between HoF models, within the convex hull $\mathcal{CH}_{\textit{train}}$ of the $\{\text{HoF}^{\textit{train}}\}$ models used for training. Given a model $\text{HoF}_{k}^{\textit{test}}$ excluded during training, we randomly sample 50 points in the model latent space close to the defining $(I_{\textit{thr}},s_{\textit{thr}})$ features of $\text{HoF}_{k}^{\textit{test}}$ (Figure˜11 illustrates the small circle around $\text{HoF}_{k}^{\textit{test}}$ in which the 50 synthetic models are sampled), none of which has been experienced during training. We run NOBLE in parallel on these sampled models to generate the corresponding somatic voltage responses and compare them to the response obtained from the numerical solver with $\text{HoF}_{k}^{\textit{test}}$. Figure˜4C shows that, when interpolating between models within the latent space, NOBLE accurately captures the underlying neuronal dynamics and produces voltage responses consistent with experimental recordings. In contrast, Figure˜4B shows that the underlying parameterizations of the HoF models obtained using the multi-objective evolutionary algorithms lack a consistent structure that would allow for meaningful interpolation to discover new bio-realistic models. This difference is further illustrated in Figure˜4A, where the F-I curves generated by the PDE models deviate markedly from experimental data, while those produced by NOBLE remain biophysically meaningful and closely aligned with observed neuronal behavior.
@@ 4.4 Ensemble Predictions
Figure 5: A) Distributions of somatic voltage traces across HoF models for current injections of $0.1\mathrm{nA}$ (top) and $0.11\mathrm{nA}$ (bottom). B) Relative errors of ensemble predictions from PDE simulations and NOBLE models on $\{\text{HoF}^{\textit{train}}\}$ compared to experimental recordings across the four key features.
@@ 4.5 Choice of Biologically-Informed Latent Embeddings
Feature embeddings are central to NOBLE as they enable generalization across and in-between biological neuron models. Here, we have chosen $I_{thr}$ and $s_{thr}$ for their strong biological interpretability, as a natural 2D representation of firing and non-firing dynamics. To quantify the contribution of embeddings, we conduct an ablation study evaluating NOBLE with lower-dimensional embeddings, as well as without embeddings. We also consider a higher-dimensional embedding that includes AHP_depth, selected for its large variation across intracell HoF models and its low correlation with $I_{thr}$ and $s_{thr}$. Results are reported in Table˜1 in terms of the relative L2 error of predicted voltage traces and the four key electrophysiological features on the test set with $\{\text{HoF}^{\textit{train}}\}$ models.
@@ 4.6 Feature Specific Physics-Informed Fine-Tuning of NOBLE
To preserve overall neural dynamics while improving accuracy on a single specific feature, NOBLE can be fine-tuned with a weighted composite loss $\mathcal{L}(\lambda)=\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{F}$, where $\mathcal{L}_{F}$ penalizes deviations in feature $F$, and $\lambda$ controls its influence. We illustrate this by fine-tuning NOBLE to improve sag_amplitude accuracy, a feature reflecting the hyperpolarization-activated cation channel (Ih). In the human cortex, Ih expression varies with cortical depth, making sag_amplitude a relevant physiological marker. We fine-tune on 19,600 subthreshold stimuli with negative amplitudes. Even without a feature-specific loss, fine-tuning reduces the L2 feature error from 70% to 19.2%, and incorporating $\mathcal{L}_{F}$ further lowers it to 9.6% while preserving overall signal fidelity. These results show that NOBLE can be selectively refined to prioritize biophysical features of interest without compromising overall performance. Further details are provided in Section˜D.7.
@@ C.3 The Fourier Neural Operator Architecture
where $\mathcal{P}$, $\mathcal{Q}$ are the pointwise neural networks that encode the lower dimension function into a higher-dimensional space and vice versa. The model stacks $L$ layers of $\sigma(W_{l}+\mathcal{K}_{l}+b_{l})$ where $W_{l}$ are pointwise linear operators (matrices), $\mathcal{K}_{l}$ are integral kernel operators, $b_{l}$ are bias terms, and $\sigma$ are fixed activation functions. The parameters $\theta$ consists of all the parameters in $\mathcal{P},\mathcal{Q},W_{l},\mathcal{K}_{l},b_{l}$. Kossaifi et al. [55] maintain a comprehensive open-source PyTorch library for learning neural operators, which serves as the foundation for our implementation. Prior knowledge of the relevant physics laws and differential equations can also be incorporated as additional loss terms during training, to supplement or replace reference data, as done with physics-informed neural operators [57, 58, 59].
A variety of neural operators have been proposed, such as the Fourier Neural Operator (FNO) [29], and successfully applied to a wide range of problems [60, 61, 62]. A FNO is a neural operator using Fourier integral operator layers, which are defined via
@@ D.4 Neighborhood Considered in Interpolation Experiments
Figure 11: Latent representations in normalized $(I_{\textit{thr}},s_{\textit{thr}})$-space of $\{\text{HoF}^{\textit{train}}\}$ (black dots) and $\{\text{HoF}^{\textit{test}}\}$ (red crosses) models. The latter lie in the convex hull $\mathcal{CH}_{\textit{train}}$ of the $\{\text{HoF}^{\textit{train}}\}$ models. In the interpolation experiment of Section˜4.3, we sample 50 models in a small neighborhood around a given model $\text{HoF}_{k}^{\textit{test}}$. We plot the boundary of this neighborhood (the blue circle) for an example of model $\text{HoF}_{k}^{\textit{test}}$. Note that this neighborhood does not contain any model used for training, and in particular, none of the 50 new models sampled in that neighborhood have been used for training.

### v4

@@ 1 Introduction
The challenges of scalability and the computational cost associated with traditional numerical modeling approaches, such as numerical integrators and evolutionary optimization algorithms, have led the scientific community to explore the use of machine learning to accelerate simulations by learning underlying relationships between variables directly from experimental and synthetic data. While neural networks have been used successfully for many applications, they learn mappings between finite-dimensional vectors, which can limit their ability to model physical phenomena that are better described using functions in infinite-dimensional spaces and functional relationships between them [25]. As a result, neural networks can overfit to the training discretization and suffer from limited out-of-distribution capabilities. Neural operators [25, 26, 27] are a principled way to generalize neural networks to learn operators mapping functions to functions, with a universal operator approximation property [28]. A variety of neural operators have been proposed, such as the Fourier Neural Operator (FNO) [29, 30].
@@ 3.3 Neural Operators for Neuron Dynamics Simulation
We choose to use neural operators as they offer clear advantages for modeling complex dynamics (Section˜C.3). Among neural operators, the Fourier Neural Operator (FNO) [29, 30] is very efficient as it leverages fast Fourier transforms on equidistant grids, which aligns naturally with our setting where both experimental recordings and PDE simulations are sampled at constant timesteps. FNOs provide a principled and efficient framework for modeling neuronal dynamics by learning mappings from input currents to voltage responses across a broad family of neuron models and current injections. Unlike conventional neural networks that operate on vector inputs and outputs of fixed sizes, the FNO learns operators, that is, mappings between functions. By operating in the frequency domain, the FNO efficiently captures global, nonlinear, and high-frequency components of voltage responses. These properties allow the model to generalize across different temporal resolutions, input currents, and neuron types, enabling the accurate simulation of unseen configurations without retraining.
@@ 4.1 Experimental Setup
For evaluating NOBLE, we focus on the PVALB neuron example introduced in Section˜2, and further assess the framework’s generality using a VIP neuron. In the PVALB setting, NOBLE receives as input the applied current injection $I$, together with stacked embeddings of $I$ (with $K=9$ different frequencies) and of the normalized model features $(I_{\textit{thr}},s_{\textit{thr}})$ associated to a neuron model $\text{HoF}_{\ell}$ (with $K=1$ frequency). NOBLE then outputs a corresponding somatic voltage response. The current injections are square DC steps with an activation duration of $400\mathrm{ms}$, consistent across all stimuli used for training and testing. We have access to 60 HoF models, where 50 are used during training, $\{\text{HoF}^{\textit{train}}\}$, and the remaining 10 HoF models $\{\text{HoF}^{\textit{test}}\}$ are used for testing. Figure˜10 displays the latent representations in normalized $(I_{\textit{thr}},s_{\textit{thr}})$-space of these HoF models. For more details on the data generation, see Section˜D.1.
@@ 4.2 Testing on HoF Models Included in the Training Set
Figure 3: A) F–I curves from experimental recordings, PDE simulations, and NOBLE predictions on $\{\text{HoF}^{\textit{train}}\}$. For one HoF model, B) compares experimental voltage responses with PDE simulations at a current injection of $0.1\mathrm{nA}$ (top) and $0.11\mathrm{nA}$ (bottom), and C) compares the corresponding PDE simulations with the NOBLE predictions for the same HoF model.
@@ 4.3 Interpolating Between Models
Figure 4: A) F-I curves from experimental recordings, PDE simulations, and NOBLE predictions on 50 interpolated HoF models. B) Experimental response vs. numerical simulation after interpolating in between PDE parameterizations. C) Numerical simulation of $\text{HoF}_{k}^{\textit{test}}$ vs. distribution of $50$ NOBLE predictions after interpolating within the model latent space near the $(I_{\textit{thr}},s_{\textit{thr}})$ features of $\text{HoF}_{k}^{\textit{test}}$.
We now test NOBLE’s ability to interpolate between HoF models within the convex hull $\mathcal{CH}_{\textit{train}}$ of the $\{\text{HoF}^{\textit{train}}\}$ models used for training. Consider a PDE model $\text{HoF}_{k}^{\textit{test}}$ excluded during training.
We first randomly sample 50 unseen synthetic models from a local neighborhood in the model latent space near the defining $(I_{\textit{thr}},s_{\textit{thr}})$ features of $\text{HoF}_{k}^{\textit{test}}$ (Figure˜11 illustrates this neighborhood). Figure˜4C shows that NOBLE accurately captures neuronal dynamics when interpolating within the latent space, as it produces a distribution of voltage responses consistent with those of the previously unseen $\text{HoF}_{k}^{\textit{test}}$. In addition, Figure˜4A shows that the F-I curves generated by NOBLE remain biophysically meaningful and closely aligned with experimentally observed neuronal behavior.
On the other hand, the parameterizations of the HoF models obtained using multi-objective evolutionary algorithms lack a consistent structure that would enable meaningful interpolation to discover new bio-realistic models. This is illustrated in Figure˜4B, where the prediction made by interpolating in between PDE parameterizations deviates significantly from experimental data. This can also be observed in Figure˜4A, where the F-I curves obtained by numerically solving PDE models constructed from slightly perturbed parameterizations of $\text{HoF}_{k}^{\textit{test}}$ deviate markedly from experimental data.
@@ 4.4 Ensemble Predictions
Figure 5: A) Distributions of somatic voltage traces across HoF and synthetic models for current injections of $0.1\mathrm{nA}$ (top) and $0.11\mathrm{nA}$ (bottom). B) Relative errors of ensemble predictions from PDE simulations and NOBLE models on $\{\text{HoF}^{\textit{train}}\}$ compared to experimental data across key features.
@@ 4.5 Choice of Biologically-Informed Latent Embeddings
Feature embeddings are central to NOBLE as they enable generalization across and in-between biological neuron models. Here, we have chosen $I_{thr}$ and $s_{thr}$ for their strong biological interpretability, as a natural 2D representation of firing and non-firing dynamics. To quantify the importance of these embeddings, we conduct an ablation study evaluating NOBLE with lower-dimensional embeddings, as well as without embeddings. We also consider a higher-dimensional embedding that includes AHP_depth, selected for its large variation across intracell HoF models and its low correlation with $I_{thr}$ and $s_{thr}$. Results are reported in Table˜1 in terms of the relative L2 error of predicted voltage traces and the four key electrophysiological features on the test set with $\{\text{HoF}^{\textit{train}}\}$ models.
@@ 4.6 Feature Specific Physics-Informed Fine-Tuning of NOBLE
To preserve overall neural dynamics while improving accuracy on a single specific feature, NOBLE can be fine-tuned with a weighted composite loss $\mathcal{L}(\lambda)=\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{F}$, where $\mathcal{L}_{F}$ penalizes deviations in feature $F$, and $\lambda$ controls its influence. We illustrate this by fine-tuning NOBLE to improve sag_amplitude accuracy, a feature reflecting the hyperpolarization-activated cation channel (Ih). In the human cortex, Ih expression varies with cortical depth, making sag_amplitude a relevant physiological marker. We fine-tune on 19,600 subthreshold stimuli with negative amplitudes. Even without a feature-specific loss, fine-tuning reduces the L2 feature error from 70% to 19.2%, and incorporating the feature loss $\mathcal{L}_{F}$ further lowers it to 9.6% while preserving overall signal fidelity. These results show that NOBLE can be selectively refined to prioritize biophysical features of interest without compromising overall performance. Further details are provided in Section˜D.7.
@@ C.3 The Fourier Neural Operator Architecture
where $\mathcal{P}$, $\mathcal{Q}$ are the pointwise neural networks that encode the lower dimension function into a higher-dimensional space and vice versa. The model stacks $L$ layers of $\sigma(W_{l}+\mathcal{K}_{l}+b_{l})$ where $W_{l}$ are pointwise linear operators (matrices), $\mathcal{K}_{l}$ are integral kernel operators, $b_{l}$ are bias terms, and $\sigma$ are fixed activation functions. The parameters $\theta$ consists of all the parameters in $\mathcal{P},\mathcal{Q},W_{l},\mathcal{K}_{l}$ and $b_{l}$. Kossaifi et al. [56] maintain a comprehensive open-source PyTorch library for learning neural operators, which serves as the foundation for our implementation. Prior knowledge of the relevant physics laws and differential equations can also be incorporated as additional loss terms during training, to supplement or replace reference data, as done with physics-informed neural operators [58, 59, 60].
A variety of neural operators have been proposed, such as the Fourier Neural Operator (FNO) [29, 30], and successfully applied to a wide range of problems [61, 62, 63]. A FNO is a neural operator using Fourier integral operator layers, which are defined via
@@ D.4 Neighborhood Considered in Interpolation Experiments
Figure 11: Latent representations in normalized $(I_{\textit{thr}},s_{\textit{thr}})$-space of $\{\text{HoF}^{\textit{train}}\}$ (black dots) and $\{\text{HoF}^{\textit{test}}\}$ (red crosses) models. The latter lie in the convex hull $\mathcal{CH}_{\textit{train}}$ of the $\{\text{HoF}^{\textit{train}}\}$ models. In the interpolation experiment of Section˜4.3, we construct a small neighborhood around a given $\text{HoF}_{k}^{\textit{test}}$ that defines a region of latent space not encountered during training, and sample 50 unseen models from this neighborhood. The boundary of this neighborhood is shown (blue circle) for an example $\text{HoF}_{k}^{\textit{test}}$ model.